OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim?

OpenStudy (anonymous):

let d=dad, let j=jim d=4j 2j+20=d sub. 2j+20=4j 20=2j j=10. Jim is 1o years old. I think this is the answer, but i might be wrong with setting up the second equation..

OpenStudy (anonymous):

i meant 10 not 1o sorry lol

OpenStudy (anonymous):

Let son' s age today is x, therefore father's age today will be 4x In 10 years father is 10 years older, therefore his age will be 4x+10 and Jim will be 10 years older as well. Hence Jim's age will be x+10, From the question Dad will be 20 years more than twice Jim's age. With that we get an equation 2(x+10) +20=4x+10 2x+20+20=4x+10 2x=30 X=15, 4x=60, Those are the ages of Jim and his Father.